OpenStudy (anonymous):

The phone company Splint has a monthly cellular plan where a customer pays a flat monthly fee and then a certain amount of money per minute used on the phone. If a customer uses 310 minutes, the monthly cost will be $145.5. If the customer uses 570 minutes, the monthly cost will be $236.5. A) Find an equation in the form where is the number of monthly minutes used and is the total monthly of the Splint plan. ? Do no use any commas in your answer. B) Use your equation to find the total monthly cost if 698 minutes are used. Answer: If 698 minutes are used, the total cost will be ?

OpenStudy (gw2011):

Let x=Amount per minute Let y=Flat fee 570x+y=236.50 310x+y=145.50 Subtract this equation from the above one to get: 260x=91 x=91/260 x=$0.35 per minute Substitute this value into one of the equations to find y: 310x+y=145.50 310(0.35)+y=145.50 108.50+y=145.50 y=$37 Flat fee TC=37+0.35x TC=37+0.35(698) TC=37+244.30 TC=$281.30 This is the total cost for 698 minutes